Solve the equation
(y - 7) - (y + 6) =9y

Found 2 solutions by checkley71, fractalier:
Answer by checkley71(8403)

Y-7-Y-6=9Y OR 9Y=-13 OR Y=-13/9

Answer by fractalier(6550)

Okay from
(y - 7) - (y + 6) = 9y
we get
y - 7 - y - 6 = 9y
-13 = 9y
y = -13/9